# Stability and Decay for the 2D Anisotropic Navier-Stokes Equations with Fractional Horizontal Dissipation on $\mathbb{R}^{2}$

### v1

@@ 4.6 The decay of $\partial_{1}u_{1}$
$\|\partial_{1}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.\vskip-5.69046pt$
For the linear term involving initial data $\partial_{1}u_{0,1}$, we can obtain the decay estimate through Lemma 2.7:
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,1}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\left(\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|\partial_{1}u_{0,1}\|_{L^{2}}\right)$
@@ 4.7 The decay of $\partial_{1}u_{2}$
$\|\partial_{1}u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+2}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\partial_{1}u_{0,2}\|_{L^{2}}\big)$
@@
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau$
@@
Then using the standard bootstrap argument, we can conclude that $T=+\infty$ and derive the conclusions in Theorem 1.5.
5.1 Estimate for $\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$
We splitting $\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}$ into two parts according to the equivalence ∥Λ_1^-σ u(t)∥_H^k ∼∥Λ_1^-σ u(t)∥_L^2 + ∑_m=1^2∥Λ_1^-σ ∂_m^k u∥_L^2, where the term $\sum_{m=1}^{2}\|\Lambda_{1}^{-\sigma}\partial_{m}^{k}u\|_{L^{2}}$ does not require coupling with decay estimates and has been completed in Subsection 4.2.
where
$\int_{0}^{t}M_{0}(\tau)d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
@@ 5.2 Weighted estimate of the pressure
The next step is to establish decay estimates for the solution. A major difficulty arises from the introduction of weighted norms: the Leray projection operator $\mathbb{P}$ is no longer directly bounded, since it involves Riesz transforms in a weighted setting. In particular, the weight $x_{2}]^{\eta}$ fails to be an $A_{2}$ when $\eta\geq\frac{1}{2}$, and therefore standard weighted boundedness results for Riesz transforms do not apply. Remarkably, the Riesz transform associated with the pressure term can still be eliminated under the weight $x_{2}]^{\eta}$. This is made possible by the special structural properties of the pressure, which allow us to bypass the lack of $A_{2}$ control. The precise formulation of this observation is given in the following lemma.
@@
Since $\frac{1}{2}<\eta<\frac{3}{2}$, we can use Lemma 2.13 and 2.10 to get
@@ 5.4 The decay of $u_{2}$ with weight
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}+\|[x_{2}]^{\gamma}u_{0,2}\|_{L^{2}}\big)$
@@ 5.5 The decay of $\partial_{2}u_{1}$ with weight
Similarly, for $III_{3}^{w}$, by applying Lemma 2.7 and Hölder’s inequality, we have
@@
By interpolation we have
@@ 5.7 The decay of $\partial_{1}u_{2}$ with weight
Applying Lemma 2.7 to (5.47) and substituting the resulting estimate together with (5.48) into (5.48), we obtain
@@ Appendix B The decay of $u_{2}$ for $0<s<\frac{3}{4}$
$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{0,2}\|_{L^{2}}\big)$
@@
$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{1}{2}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\|_{L^{2}}\Big)d\tau.$
@@
$\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{1}{2}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\|_{L^{2}}\Big)d\tau$

### v2

@@ 4.5 The decay of $\partial_{1}u_{1}$
$\|\partial_{1}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{1,0}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.\vskip-5.69046pt$
For the linear term involving initial data $\partial_{1}u_{1,0}$, we can obtain the decay estimate through Lemma 2.7:
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{1,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\left(\|\Lambda_{1}^{-\sigma}u_{1,0}\|_{L^{2}}+\|\partial_{1}u_{1,0}\|_{L^{2}}\right)$
@@ 4.6 The decay of $\partial_{1}u_{2}$
$\|\partial_{1}u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{2,0}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u)(\tau)\|_{L^{2}}d\tau.$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}u_{2,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+2}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|\partial_{1}u_{2,0}\|_{L^{2}}\big)$
@@
$\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}^{2}\Delta^{-1}\nabla\times\mbox{div}\ (u\otimes u)(\tau)\|_{L^{2}}d\tau$
@@
4.7 Estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$
In this subsection, we show that, under the smallness assumption on $\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}$, the decay rates of $u$ and its derivatives are sufficient to guarantee the propagation of $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$.
$:=$ $M_{0}.$
Integrating in time on equation (4.38), we obtain
By Lemma 2.6,
Using the divergence-free condition of $u$, we have
It can be verified that
$s<\tfrac{5}{12}+\sigma<\tfrac{-18\sigma^{2}+25\sigma+3}{12}.$
As a consequence, we obtain
$\int_{0}^{t}M_{0}(\tau)\,d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
@@
Then using the standard bootstrap argument, we can conclude that $T=+\infty$ and derive the conclusions in Theorem 1.5. At the end, based on the decay rates of $u$ and its derivatives, we also prove the propagation of the negative horizontal derivative of $u$.
@@ 5.1 Weighted estimate of the pressure
The energy estimate of $\|u\|_{H^{k}}$ is the same as the process in subsection 4.1. So the next step is to establish decay estimates for the solution. A major difficulty arises from the introduction of weighted norms: the Leray projection operator $\mathbb{P}$ is no longer directly bounded, since it involves Riesz transforms in a weighted setting. In particular, the weight $x_{2}]^{\eta}$ fails to be an $A_{2}$ when $\eta\geq\frac{1}{2}$, and therefore standard weighted boundedness results for Riesz transforms do not apply. Remarkably, the Riesz transform associated with the pressure term can still be eliminated under the weight $x_{2}]^{\eta}$. This is made possible by the special structural properties of the pressure, which allow us to bypass the lack of $A_{2}$ control. The precise formulation of this observation is given in the following lemma.
@@
Since $\frac{1}{2}<\eta<\frac{3}{2}$, we can use Lemma 2.13 and Lemma 2.10 to get
@@ 5.3 The decay of $u_{2}$ with weight
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{1}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}[x_{2}]^{\gamma}\psi_{0}\|_{L^{2}}+\|[x_{2}]^{\gamma}u_{2,0}\|_{L^{2}}\big)$
@@ 5.4 The decay of $\partial_{2}u_{1}$ with weight
Similarly, for $III_{3}^{w}$, by applying Lemma 2.7 and Hölder inequality, we have
@@
Applying interpolation to the second term in the integrand, we obtain
@@ 5.6 The decay of $\partial_{1}u_{2}$ with weight
Applying Lemma 2.7 to (5.46) and substituting the resulting estimate in (5.47) into (5.45), we obtain
@@
5.7 Estimate for $\|\Lambda_{1}^{-\sigma}u\|_{L^{2}}$
In this subsection, under the smallness assumption on $\|\Lambda_{1}^{-\sigma}u_{0}\|_{L^{2}}$, we prove that the decay rates of $u$ and its derivatives imply that $\|\Lambda_{1}^{-\sigma}u(t)\|_{L^{2}}$ is finite for all time.
where
$\int_{0}^{t}M_{0}(\tau)d\tau\leq CC_{0}^{3}\varepsilon^{3}.$
@@ Appendix B The decay of $u_{2}$ for $0<s<\frac{3}{4}$
$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{2,0}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}u_{2,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{2,0}\|_{L^{2}}\big)$
@@
$\leq$ $C\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}\Big[(\Lambda_{1}^{\frac{1}{2}}u_{1})u\Big]\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}\Big[u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\Big]\|_{L^{2}}\Big)d\tau.$
@@
$\int_{0}^{t-1}(t-\tau)^{-\frac{\sigma+1}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}\Big[(\Lambda_{1}^{\frac{1}{2}}u_{1})u\Big]\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}\Big[u_{1}(\Lambda_{1}^{\frac{1}{2}}u)\Big]\|_{L^{2}}\Big)d\tau$